# Nonperturbative cavity quantum electrodynamics: is the Jaynes-Cummings model still relevant?

### v1

@@
journal: opticajournalarticletype: Research Article{abstract*}
@@ 2.1 Introducing the Jaynes-Cummings model
where for each $n>0$ the eigenvalues of ${}^{\text{th}}$ block, $\omega_{n,\pm}$, describe dressed states with $n$ excitations. The saturation of the TLS makes the system nonlinear, leading to a $n$-dependent intra-doublet splitting which at resonance reads [17, 18]
@@ 2.2 Why we need to go beyond the JCM: three experimental examples
System Cavity $\omega_{c}$ Atom $\omega_{eg}$ Coupling $\Omega_{\rm R}$
$\zeta$
Superconducting circuits [19]
35.2 GHz
23.9 GHz
35.2 GHz
6
Molecular plasmonic cavities [21]
452 THz
452 THz
73 THz
0.03
Graphene quantum dots [20]
25 THz
3.8 THz
49 THz
101
@@ 3.1 Dipolar approximation
To start, the JCM considers the light-matter interaction in the dipolar approximation. Formally this is obtained by expressing the full non-relativistic QED Hamiltonian in the so-called Poincaré gauge [42]: truncating its multipolar expansion to the lowest order, one obtains the dipole gauge Hamiltonian [43]. In simpler terms, when the electromagnetic field does not vary too much on the length scales of the TLS spatial extension, it can be considered spatially uniform. As a consequence, the light-matter interaction Hamiltonian can be derived by considering the energy of a dipole $\hat{\mathbf{d}}$ in a uniform electric field $\hat{\mathbf{E}}$
$\hat{H}_{I}\propto\hat{\mathbf{d}}\cdot\hat{\mathbf{E}}.$ (9)
@@ 3.2 Modelling the emitter as a TLS
While transitions in spin doublets can be exactly modeled as TLS, most quantum emitters are implemented with electronic transitions. The number of trapped electronic states can be substantially larger than two, and the possibility of focusing only on a single, discrete transition, between the ground state ($|g\rangle$) and a single excited state ($|e\rangle$) rests on the assumption that the coupling with all the other parts of the spectrum can be neglected. The evolution of the system can thus be considered to span only the two-dimensional Hilbert space [|$\left[\ket{g},\ket{e}\right]$. Using the Pauli matrices we can then describe any operator in the two-level subspace, for instance, the atomic dipole operator becomes
@@ 3.3 Considering a single cavity mode
Analogously to the matter degrees of freedom, the photonic resonator also hosts a complete set of modes [56], both discrete and belonging to a continuum. If the field is strongly confined in a cavity or in any artificial structure (e.g. subwavelength resonators), the energy spacing between the different electromagnetic modes is such that we can discard all of them except the one that is most resonant with the optically active transition of interest. In such a way the cavity dynamics is completely described as a harmonic oscillator, with annihilation operator $\hat{a}$, and the electric field operator is given by
$\hat{\mathbf{E}}\propto\hat{a}+\hat{a}^{{\dagger}}.$ (11)
@@ 3.4 Applying the rotating wave approximation
The rotating wave approximation (RWA) is ubiquitous in physics even if known with different names, another one being secular approximation from its use in celestial mechanics, where the errors introduced would have become observable only over centuries. In the context of CQED is implemented by the following reduction
@@ 4.4 Excited states beyond RWA: multi-photon non-linear processes
Increasing the value of $\epsilon$ opens some non-trivial avoided crossings in the spectrum, allowing the system to undergo multi-photon Rabi oscillations that might be used for the generation of multi-photons Fock states [81, 79, 82]. Here a single atom can emit simultaneously multiple photons with a single transition due to the USC. Differently from multi-photon processes arising in devices with non-linear quadrupolar light-matter coupling [83, 49], this dynamics is a USC consequence of the interplay between the linear dipole coupling and the non-linearity (or saturability) of the TLS. These effects are typically called tunneling resonances [84] since their mathematical description is identical to tunneling resonances appearing in the physics of electronic transport assisted by phonons through molecular or nanostructure quantum dots [85, 86, 87, 88]. They are currently a major reason for interest in the development of these platforms projecting also interesting new technological perspectives on the USC regime [89, 90, 91, 37, 36].
@@ 5 Gauging effective models
To better understand the problem, let’s consider the two most used gauges to study non-relativistic light-matter interactions: the Coulomb and the Power-Zienau-Woolley gauges [99, 42], where the latter is also known as multipolar or dipole gauge. Both gauges are also dubbed in the literature as the $\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}$ or velocity gauge and the $\hat{\mathbf{d}}\cdot\hat{\mathbf{E}}$ or length gauge, where $\hat{\mathbf{p}}$ and $\hat{\mathbf{d}}=q\hat{\mathbf{x}}$ are the particle momentum and electric dipole moment with charge $q$ and displacement $\hat{\mathbf{x}}$. $\hat{\mathbf{A}}$ and $\hat{\mathbf{E}}$ are the vector potential and the electric field. If we consider the transition matrix elements between two states $|j\rangle$ and $|k\rangle$ of a particle with mass $m$, position $\hat{\mathbf{x}}$, and momentum $\hat{\mathbf{p}}$ we can easily derive the equation [100]
@@
The same results can have also been obtained following a lattice gauge theory approach [103].
@@ 6.1 Bosonising the light-matter interactions
Here the index $i$ on the Pauli matrices addresses the different TLSs. We introduced the collective annihilation operator $\hat{b}=\sum_{i=1}^{N}\hat{\sigma}_{-}^{i}/\sqrt{N}$, describing a collective excitation of all the TLSs. This operator satisfies $[\hat{b},\hat{b}^{{\dagger}}]\approx 1+O(\frac{n_{x}}{N})$ in the dilute regime, that is when the number of excitations $n_{x}$ is much smaller than the total number $N$ of TLSs [110, 111].
@@
In Fig. 4(a)-(c) we show three paradigmatic examples of CQED setups well described by the bosonic approximation: intersubband polaritons [112], Landau polaritons [113] and magnonic polaritons [114]. In order to recover the technologically relevant nonlinear regime, there has been a constant effort in those systems to reduce the number of dipoles while keeping the system in the USC regime [115, 116, 5, 117]. Spectral differences at the transition between collective and single particle physics are shown in Fig. 4(d).
@@ 7 Open quantum systems: can we measure the non-empty vacuum?
This problem can be solved by not performing the white-reservoir approximation and deriving the Liouvillian in the eigenbasis of the light-matter Hamiltonian [132]. The resulting dressed master equation takes the form
@@
where the dissipator is expressed in terms of jump operators between the $k$-th and the $j$-th eigenstates of the coupled light-matter Hamiltonian, |$\outerproduct{k}{j}$ |, with the associated Bohr frequency $\omega_{jk}=\omega_{j}-\omega_{k}$. The vanishing density of states of the bath at negative frequencies then leads to $\gamma(\omega<0)=0$, solving the problem of energy conservation. As a consequence, the open dynamics must be interpreted as transitions between the true eigenstates, where light and matter are entangled together and it is not possible to simply distinguish them separately, as schematized in Fig. 5(a-b)
A problem with this formulation can nevertheless emerge when using it in the weak coupling regime, due to the degenerate nature of the spectrum in such a regime. This in turn led to the derivation of generalized master equations  [130] explicitly taking these degeneracies into account and leading to correct results for all values of the light-matter coupling.
@@
which gives $\langle\mathrm{G}_{\mathrm{R}}|\hat{E}^{-}\hat{E}^{+}|\mathrm{G}_{\mathrm{G}}\rangle=0$. Moreover, the form of the electric field operator becomes gauge dependent [119, 136, 137], but leaving the photon rate in Eq. 25 gauge-invariant.
Aside from the correct formalism to adopt, these issues make us to reflect on the meaning of photons in the presence of matter, a point that was raised already a while ago [138]. Virtual photons arising from the USC to matter cannot be simply interpreted as quanta of the transverse electric field oscillations, as commonly done in the free space case. Their physical meaning depends on the gauge that we adopt to describe the interaction with matter and it must be handled with care [32, 94, 139, 140].
@@ 8.2 The gRWA and the polaronic JCM
The origin of the USC light-matter decoupling in single mode systems can be traced back to the existence of a hidden approximation: the generalized rotating-wave approximation (gRWA). As it was first reported in Ref. [146] and then developed in Ref. [147], it is possible to recover the simple physics of the JCM when the coupling strength places the system deep into the USC regime, where $\Omega_{\rm R}/\omega_{c}\gg 1$. As a consequence, and apparently contradicting what we reported in the previous Sections, the ground state of the system appears as a trivial empty vacuum.
The gRWA is not the standard RWA that leads to the JCM because it first requires transforming the Rabi Hamiltonian in Eq. (13) in a new basis, often called polaron frame. The coordinate transformation is implemented by the same unitary transformation in Eq. (19) which was used to implement the minimal substitution in the Coulomb gauge Hamiltonian directly in the truncated two-level subspace. In this context, unrelated to gauge choices, this transformation is known as polaron transformation. The polaron frame Rabi Hamiltonian (which coincides with the Coulomb gauge TLS Hamiltonian presented at the end of Sec. 5) is then given by
@@
where $f$ is a complicated polynomial function [146]. This expression is still mathematically very complicated but has the advantage of clearly showing that also the Rabi model is mostly built over the concept of absorption/emission. The exponential coefficient in front $e^{-2\Omega_{\rm R}^{2}/\omega_{c}^{2}}$ emerges also as a feature of the polaron transformation and explicitly accounts for the light-matter decoupling in the infinite coupling limit $\Omega_{\rm R}\rightarrow\infty$. It also explains why the gRWA is applicable in such a limit, suppressing the polaron light-matter interaction term which is then treated in perturbation theory, similarly to what is introduced in Sec. 2.
@@
where $|\alpha\rangle$ is a photon’s coherent state with amplitude $\alpha=\Omega_{\rm R}/\omega_{c}$ and $|\leftarrow/\rightarrow\rangle=(|e\rangle\pm|g\rangle)/\sqrt{2}$. We then notice that all the notions of vacuum and virtual photons are only relative to the specific considered frame (or Hilbert space basis). Moreover, we see that the Rabi vacuum is not only non-empty but is also highly entangled since for $\Omega_{\rm R}\gg\omega_{c}$ it is a cat state [1]. In the polaron basis, the ground state instead becomes quite trivial and loses all the entanglement (which is also a relative property). This observation stimulated the idea that in light-matter problems one can always find a disentangling transformation that strongly simplifies the description. In particular, this was explored in multimodal light-matter systems, where the polaron transformation is generalized to treat the USC regime on the basis of minimizing the entanglement between matter and light. In Ref. [149, 150] is shown that a generalized polaron transformation can be used to extract semi-analytical approximated solutions of the multi-mode USC systems.

### v2

@@
journal: opticajournalarticletype: Research Article
{abstract*}
@@ 2.1 Introducing the Jaynes-Cummings model
where for each $n>0$ the eigenvalues of ${}^{\text{th}}$ block, $\omega_{n,\pm}$, describe dressed states with $n$ excitations. The JCM spectrum is then immediately given by $\omega_{n,\pm}=-\omega_{eg}/2+\omega_{c}n-\Delta/2\pm\sqrt{\Delta^{2}/4+\Omega_{\text{R}}^{2}n}$, where we have introduced the cavity-TLS detuning $\Delta=\omega_{c}-\omega_{eg}$. The saturation of the TLS makes the system nonlinear, leading to a $n$-dependent intra-doublet splitting which at resonance, $\Delta=0$, reads [17, 18]
@@
The JCM nonlinearity is at the base of many important effects in quantum optics, among all it is worth mentioning the photon blockade [19, 20, 21, 22], which is a central object in the developments of quantum technology [23, 24].
@@ 2.2 Why we need to go beyond the JCM: three experimental examples
System Cavity $\omega_{c}$ Atom $\omega_{eg}$ Coupling $\Omega_{\rm R}$ $\zeta$
Superconducting circuits [25] 35.2 GHz 23.9 GHz 35.2 GHz 6
Molecular plasmonic cavities [27] 452 THz 452 THz 73 THz 0.03
Graphene quantum dots [26] 25 THz 3.8 THz 49 THz 101
@@ 3.1 Dipolar approximation
To start, the JCM considers the light-matter interaction in the dipolar approximation. Formally this is obtained by expressing the full non-relativistic QED Hamiltonian in the so-called Poincaré gauge [48]: truncating its multipolar expansion to the lowest order, one obtains the dipole gauge Hamiltonian [49]. In simpler terms, when the electromagnetic field does not vary too much on the length scales of the TLS spatial extension, it can be considered spatially uniform. As a consequence, the light-matter interaction Hamiltonian can be derived by considering the energy of an electric dipole $\hat{\mathbf{d}}$ in a uniform electric field $\hat{\mathbf{E}}$ or a magnetic dipole $\hat{\mathbf{m}}$ in a uniform magnetic field $\hat{\mathbf{B}}$
$\hat{H}_{I}\propto\hat{\mathbf{d}}\cdot\hat{\mathbf{E}}\quad{\rm or}\quad\hat{\mathbf{m}}\cdot\hat{\mathbf{B}}.$ (9)
While standard cavity QED discussed in textbook is typically due to the electric dipole coupling [50, 51, 18], it is worth stressing that a magnetic coupling leads to the very same phenomenology either through the Zeeman term [52, 53, 54, 55, 56, 57] or through orbital magnetism [58, 59, 60].
@@ 3.2 Modelling the emitter as a TLS
While transitions in spin doublets can be exactly modeled as TLS, most quantum emitters are implemented with electronic transitions. The number of trapped electronic states can be substantially larger than two, and the possibility of focusing only on a single, discrete transition, between the ground state ($|g\rangle$) and a single excited state ($|e\rangle$) rests on the assumption that the coupling with all the other parts of the spectrum can be neglected. The evolution of the system can thus be considered to span only the two-dimensional Hilbert space [|$\left[\ket{g},\ket{e}\right]$. Using the Pauli matrices we can then describe any operator in the two-level subspace, for instance, the atomic electric dipole operator becomes
@@ 3.3 Considering a single cavity mode
Analogously to the matter degrees of freedom, the photonic resonator also hosts a complete set of modes [73], both discrete and belonging to a continuum. If the field is strongly confined in a cavity or in any artificial structure (e.g. subwavelength resonators), the energy spacing between the different electromagnetic modes is such that we can discard all of them except the one that is most resonant with the optically active transition of interest. In such a way the cavity dynamics is completely described as a harmonic oscillator, with annihilation operator $\hat{a}$, and the electric and magnetic field operators are given by
$\hat{\mathbf{E}}\propto\hat{a}+\hat{a}^{{\dagger}},\quad\hat{\mathbf{B}}\propto-i\left(\hat{a}-\hat{a}^{{\dagger}}\right).$ (11)
@@ 3.4 Applying the rotating wave approximation
The rotating wave approximation (RWA) is ubiquitous in physics even if known with different names, another one being secular approximation from its use in celestial mechanics, where the errors introduced would have become observable only over centuries. In the context of CQED is implemented by the following reduction (the same holds for the magnetic coupling)
@@ 4.4 Excited states beyond RWA: multi-photon non-linear processes
Increasing the value of $\epsilon$ opens some non-trivial avoided crossings in the spectrum, allowing the system to undergo multi-photon Rabi oscillations that might be used for the generation of multi-photons Fock states [97, 95, 98]. Here a single atom can emit simultaneously multiple photons with a single transition due to the USC. Differently from multi-photon processes arising in devices with non-linear quadrupolar light-matter coupling [99, 66], this dynamics is a USC consequence of the interplay between the linear dipole coupling and the non-linearity (or saturability) of the TLS. These effects are typically called tunneling resonances [100] since their mathematical description is identical to tunneling resonances appearing in the physics of electronic transport assisted by phonons through molecular or nanostructure quantum dots [101, 102, 103, 104]. They are currently a major reason of interest in the development of these platforms projecting also new technological perspectives on the USC regime, in terms of new devices for parametric up/down conversion and multi-photon Fock state preparation [105, 106, 107, 108, 43, 42].
@@ 5 Gauging effective models
To better understand the problem, let’s consider the two most used gauges to study non-relativistic light-matter interactions: the Coulomb and the Power-Zienau-Woolley gauges [116, 48], where the latter is also known as multipolar or dipole gauge. Both gauges are also dubbed in the literature as the $\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}$ or velocity gauge and the $\hat{\mathbf{d}}\cdot\hat{\mathbf{E}}$ or length gauge. Here $\hat{\mathbf{p}}$ and $\hat{\mathbf{d}}=q\hat{\mathbf{x}}$ are the particle momentum and electric dipole moment with charge $q$ and displacement $\hat{\mathbf{x}}$. $\hat{\mathbf{A}}$ and $\hat{\mathbf{E}}$ are the vector potential and the electric field. Notice that in many text-books is conventionally adopted the dipole gauge in the derivation of the JCM, motivating our choice in Eq. (9) instead of reporting the dipole coupling in terms of Coulomb gauge coupling. If we consider the transition matrix elements between two states $|j\rangle$ and $|k\rangle$ of a particle with mass $m$, position $\hat{\mathbf{x}}$, and momentum $\hat{\mathbf{p}}$ we can easily derive the equation [117]
@@
While the results are completely equivalent to what was described above, this approach represents the basis of lattice gauge theories [120], becoming another interesting example of how cavity QED can be a useful playground to experiment with the most complex concepts of modern physics. It is worth mentioning that, as much as in other branches of physics [121, 122], the discussion about gauge invariance has risen an intense debate in the community, which is still on going [123, 124]. Moreover, while here we focused mainly on Coulomb and dipole gauge, other gauge choices may be also convenient, dependently from the specific problem the one needs to discuss [118].
We conclude this section by noticing that both dipole and Coulomb gauge returns a Hamiltonian that contains the counter-rotating terms in the light-matter coupling. However it was shown in Refs. [125, 111] that is possible to implement an intermediate gauge transformation which exactly cancels the counter-rotating terms. In this representation the RWA is no longer an approximation and the vacuum is always empty. While we will come back on the relative meaning of the quantum vacuum in the last section, it is important to understand that, under the standard assumptions given above and described in Ref. [68], this representation pays the price of being not fully compatible with the TLA as much as the Coulomb gauge [115].
@@ 6.1 Bosonising the light-matter interactions
Here the index $i$ on the Pauli matrices addresses the different TLSs. We introduced the collective annihilation operator $\hat{b}=\sum_{i=1}^{N}\hat{\sigma}_{-}^{i}/\sqrt{N}$, describing a collective excitation of all the TLSs. This operator satisfies $[\hat{b},\hat{b}^{{\dagger}}]\approx 1+O(\frac{n_{x}}{N})$ in the dilute regime, that is when the number of excitations $n_{x}$ is much smaller than the total number $N$ of TLSs [132, 133]. It is worth noticing that in this same limit, the equilibrium Dicke model is exactly solvable [134, 135, 136], making it a paradigmatic example to discuss basics many-body effects in light-matter interactions.
@@
In Fig. 4(a)-(c) we show three paradigmatic examples of CQED setups well described by the bosonic approximation: intersubband polaritons [137], Landau polaritons [138] and magnonic polaritons [139]. In order to recover the technologically relevant nonlinear regime [140], there has been a constant effort in those systems to reduce the number of dipoles while keeping the system in the USC regime [141, 142, 5, 143]. Spectral differences at the transition between collective and single particle physics are shown in Fig. 4(d). It is important to point out that Bosonized excitations with non-negligible non-linearities are also found in phonon-polaritons THz materials, which is another great area of interest regarding the USC impact on matter [144, 145, 118, 119, 146, 147].
@@ 7 Open quantum systems: can we measure the non-empty vacuum?
This problem can be solved by not performing the white-reservoir approximation and deriving the Liouvillian in the eigenbasis of the light-matter Hamiltonian [157, 162]. The resulting dressed master equation takes the form
@@
where the dissipator is expressed in terms of jump operators between the $k$-th and the $j$-th eigenstates of the coupled light-matter Hamiltonian, |$\outerproduct{k}{j}$ |, with the associated Bohr frequency $\omega_{jk}=\omega_{j}-\omega_{k}$. Consistently with the secular approximation (or just RWA) needed to obtain the master equation in a Linblad form, one needs to discard all the transitions with negative frequencies that simultaneously create an excitation in the bath, for which in the above expression we must enforce that $j>k$. The problem of energy conservation is now solved, being only an artifact of a wrong derivation. As a consequence, the open dynamics must be interpreted as transitions between the true eigenstates, where light and matter are entangled together and it is not possible to simply distinguish them separately, as schematized in Fig. 5(a-b) The above derivation is well justified at weak system-bath coupling and it can be equivalently rephrased as a vanishing density of states of the bath at negative frequencies leading to $\gamma(\omega<0)=0$.
It is worth noticing that the above approach relies on having a different bath for each transition [162]. This remains true even if the all the system’s transitions decay in the same bath, provided that their coupling to the bath is smaller than their relative distance in frequencies. In the weak coupling regime of cavity QED for instance this condition is removed and a problem with this formulation can nevertheless emerge due to the degenerate nature of the spectrum in such a regime. This in turn led to the derivation of generalized master equations  [160] explicitly taking these degeneracies into account and leading to correct results for all values of the light-matter coupling. Similar issues are partially discussed also in Refs. [163, 164, 165] and, more completely in a very extensive way, in Ref. [166].
@@
which gives $\langle\mathrm{G}_{\mathrm{R}}|\hat{E}^{-}\hat{E}^{+}|\mathrm{G}_{\mathrm{G}}\rangle=0$. Moreover, the form of the electric field operator becomes gauge dependent [149, 169, 170], but leaving the photon rate in Eq. 25 gauge-invariant. Similar analysis was carried out on pure dephasing processes, where the system experienced a loss of coherence due to stochastic fluctuations of the bare energies of the systems [162, 171].
Aside from the correct formalism to adopt, these issues make us reflect on the meaning of photons in the presence of matter, a point that was raised already a while ago [172]. Virtual photons arising from the USC to matter cannot be simply interpreted as quanta of the transverse electric field oscillations, as commonly done in the free space case. Their physical meaning depends on the gauge that we adopt to describe the interaction with matter and it must be handled with care [38, 111, 173, 174].
@@ 8.2 The gRWA and the polaronic JCM
The origin of the USC light-matter decoupling in single-mode systems can be traced back to the existence of a hidden approximation: the generalized rotating-wave approximation (gRWA). As it was first reported in Ref. [180] and then developed in Ref. [181], it is possible to recover the simple physics of the JCM when the coupling strength places the system deep into the USC regime, where $\Omega_{\rm R}/\omega_{c}\gg 1$. As a consequence, and apparently contradicting what we reported in the previous Sections, the ground state of the system appears as a trivial empty vacuum.
The gRWA is not the standard RWA that leads to the JCM because it first requires transforming the Rabi Hamiltonian in Eq. (13) in a new basis, often called polaron frame [182, 183]. The coordinate transformation is implemented by the same unitary transformation in Eq. (19) which was used to implement the minimal substitution in the Coulomb gauge Hamiltonian directly in the truncated two-level subspace. In this context, unrelated to gauge choices, this transformation is known as polaron transformation. The polaron frame Rabi Hamiltonian (which coincides with the Coulomb gauge TLS Hamiltonian presented at the end of Sec. 5) is then given by
@@
where $f$ is a complicated polynomial function [180]. This expression is still mathematically very complicated but has the advantage of clearly showing that also the Rabi model is mostly built over the concept of absorption/emission. The exponential coefficient in front $e^{-2\Omega_{\rm R}^{2}/\omega_{c}^{2}}$ can be seen as a non-perturbative Lamb-shift [185, 41], emerging also as a feature of the polaron transformation, and explicitly accounts for the light-matter decoupling in the infinite coupling limit $\Omega_{\rm R}\rightarrow\infty$. It also explains why the gRWA is applicable in such a limit, suppressing the polaron light-matter interaction term which is then treated in perturbation theory, similarly to what is introduced in Sec. 2.
@@
where $|\alpha\rangle$ is a photon’s coherent state with amplitude $\alpha=\Omega_{\rm R}/\omega_{c}$ and $|\leftarrow/\rightarrow\rangle=(|e\rangle\pm|g\rangle)/\sqrt{2}$. We then notice that all the notions of vacuum and virtual photons are only relative to the specific considered frame (or Hilbert space basis). Moreover, we see that the Rabi vacuum is not only non-empty but is also highly entangled since for $\Omega_{\rm R}\gg\omega_{c}$ it is a cat state [1]. In the polaron basis, the ground state instead becomes quite trivial and loses all the entanglement (which is also a relative property). This observation stimulated the idea that in light-matter problems one can always find a disentangling transformation that strongly simplifies the description. In particular, this was explored in multimodal light-matter systems, where the polaron transformation is generalized to treat the USC regime on the basis of minimizing the entanglement between matter and light. In Ref. [185, 186, 163] is shown that a generalized polaron transformation can be used to extract semi-analytical approximated solutions of the multi-mode USC systems.
To close this section, we notice that the existence of a disentangling transformation such as the polaron one does not immediately imply the aforementioned light-matter decoupling since that is a property of the full dynamics and not just of the groundstate. However, in the contest, for instance, of the Rabi model, it realizes an optimal basis from which the light-matter decoupling phenomenon emerges quite clearly. In this sense one should interpret the implication in a reversed way: because of the light-matter decoupling a disentanged basis offers a simpler and more natural view of the system. In this perspective, while light-matter decoupling exists on every basis, its manifestation in formalism can be very different from one representation to another. For instance, it was recently shown [118] that considering a different frame than the polaron one the light-matter decoupling can be seen as the increasing of the electron effective mass $m_{\mathrm{eff}}=m[1+2(\Omega_{\mathrm{R}}/\omega_{c})^{2}]$, giving rise to a tight localization of the electron around the potential minima.